OpenStudy (anonymous):

kent walked 3/4 of a mile on monday. on tuesday he walked 1/8 of a mile less than on monday. how far did he walk altogether?

OpenStudy (anonymous):

M=3/4 T=3/4-1/8 Total=3/4+3/4-1/8

OpenStudy (anonymous):

6/4-1/8 12/8-1/8 11/8 or 1 3/8 miles

OpenStudy (anonymous):

\[\frac{3}{4}=\frac{6}{8} \]\[\frac{6}{8}+\frac{5}{8}=\frac{11}{8}=1\frac{3}{8}=1.375 \text{ miles} \]